# NANOGrav hints for first-order confinement-deconfinement phase transition in different QCD-matter scenarios

### 04 Dec 2023 v1

@@ Abstract
For each observed frequency $f_{i}$, we employ the posteriors of $\hat{\Omega}_{\mathrm{GW}}(f_{i})$ obtained in Eq. (13) to estimate the corresponding kernel density, $\mathcal{L}_{i}(\hat{\Omega}_{\mathrm{GW}})$. Consequently, the total log-likelihood is the summation of individual likelihoods given by [53, 57, 63, 38]
@@
where $\Lambda\equiv\{\alpha,T_{*}\}$ represents the two model parameters. We utilize the dynesty [126] sampler available within the Bilby [127, 128] package to explore the parameter space.
@@
Summary and discussion. In this letter, we firstly demonstrate that the stochastic signal detected in the NANOGrav 15-year data set can be successfully attributed to an SGWB arising from the first-order cosmological confinement-deconfinement phase transition of either heavy static quarks with zero baryon chemical potential or quarks with finite baryon chemical potential, aligning with the results in our previous work [88]. The baryon chemical potential exerts negligible influence on the power spectra of GWs generated by confinement-deconfinement phase transitions. This is attributed to the close proximity of the phase transition temperatures in both the finite baryon chemical potential and zero chemical potential scenarios. Furthermore, it is worth noting that, based on the Bayesian method, our analysis has revealed that the phase transition of a pure gluon plasma cannot be regarded as the potential source of the NANOGrav signal, and the non-runaway bubble dynamics model is effectively ruled out by the NANOGrav data at a significance level of $5\sigma$. Secondly, we obtain a remarkably distinct result compared to the previous work, i.e., the bubble dynamics in the phase transition should be characterized by the Jouguet detonation rather than a non-runaway scenario. Thirdly, our results provide some insights into the applicability of holographic descriptions of the quark phase transitions, in which the models of hard and soft walls relate to the introduction of sharp and smooth cutoffs respectively. According to our analysis, the holographic description using the hard wall model for the quark phase transition is more likely to be consistent with PTA data as compared to that using the soft wall model.
In this work, we only examine the first-order confinement-deconfinement phase transitions. However, it is also worth considering other types of the first-order phase transition at the QCD scale, such as the chiral symmetry-breaking [129]. Moreover, relying solely on PTA observations might not be sufficient to distinguish the various details of different first-order phase transitions. Combining additional constraints [130] related to first-order phase transitions will help us gain further insights into the processes of cosmic phase transitions. We leave these aspects for future research.

### 23 Jan 2024 v2

@@ III Data analyses and results
For each observed frequency $f_{i}$, we employ the posteriors of $\hat{\Omega}_{\mathrm{GW}}(f_{i})$ obtained in Eq. (13) to estimate the corresponding kernel density, $\mathcal{L}_{i}(\hat{\Omega}_{\mathrm{GW}})$. Consequently, the total log-likelihood is the summation of individual log-likelihoods given by Moore and Vecchio (2021); Lamb et al. (2023); Liu et al. (2023b); Wu et al. (2023c); Jin et al. (2023); Liu et al. (2023a)
@@
where $\Lambda\equiv\{\alpha,T_{*}\}$ represents the two model parameters. We utilize the dynesty Speagle (2020) sampler available within the Bilby Ashton et al. (2019); Romero-Shaw et al. (2020) package to explore the parameter space. To evaluate the model fits, we calculate the Bayesian Information Criterion (BIC) Schwarz (1978), defined as
$\mathrm{BIC}=k\ln n-2\ln\hat{\mathcal{L}},$ (16)
where $n=66$ corresponds to the number of data points in the frequency space, $k$ is the number of parameters in the model, and $\hat{\mathcal{L}}$ is the maximum likelihood. The BIC facilitates the comparison of two competing models Kass and Raftery (1995): if the difference in BICs between the two models is $0-2$, it constitutes ‘weak’ evidence in favor of the model with the smaller BIC; a difference in BICs between 2 and 6 indicates ‘positive’ evidence; a difference in BICs between 6 and 10 signifies ‘strong’ evidence; and a difference in BICs greater than 10 indicates ‘very strong’ evidence in favor of the model with a smaller BIC.
@@
In Table 2, we present the differences in BIC among the ten QCD-like models with $\beta/H_{*}=10$. Remarkably, the $S_{1}$ and $S_{3}$ models within the Jouguet bubble scenario exhibit superior performance compared to all other models. Furthermore, there is ‘positive’ evidence supporting the preference for the $S_{3}$ model over the $S_{1}$ model in this scenario. Extending the parameter $\beta/H_{*}$ to a larger value, such as $\beta/H_{*}=15$, as shown in Table 3, the $S_{1}$ and $S_{3}$ models in the Jouguet bubble scenario maintain their superior performance over other models, with ‘very strong’ evidence favoring the $S_{3}$ model over the $S_{1}$ model. Therefore, a larger value of $\beta/H_{*}$ does not significantly alter our findings.
$S_{1}$ $S_{2}$ $S_{3}$ $S_{4}$ $S_{5}$
Jouguet 0 42.5 -3.1 41.2 63.4
non-runaway 43.0 60.2 34.0 60.1 59.2
Table 2: The difference of BIC among the ten QCD-like models with fixing $\beta/H_{*}=10$. We have set the $S_{1}$ from the Jouguet bubble case as the fiducial model.
$S_{1}$ $S_{2}$ $S_{3}$ $S_{4}$ $S_{5}$
Jouguet 0 17.5 -12.9 17.4 17.0
non-runaway 16.6 18.6 19.9 18.6 16.3
Table 3: Same as Table 2, but fixing $\beta/H_{*}=15$.
IV Summary and discussion
In this paper, we firstly demonstrate that the stochastic signal detected in the NANOGrav 15-year data set can be successfully attributed to an SGWB arising from the first-order cosmological confinement-deconfinement phase transition of either heavy static quarks with zero baryon chemical potential or quarks with finite baryon chemical potential, aligning with the results in our previous work Li et al. (2021). The baryon chemical potential exerts negligible influence on the power spectra of GWs generated by confinement-deconfinement phase transitions. This is attributed to the close proximity of the phase transition temperatures in both the finite baryon chemical potential and zero chemical potential scenarios. Furthermore, it is worth noting that, based on the Bayesian method, our analysis has revealed that the phase transition of a pure gluon plasma cannot be regarded as the potential source of the NANOGrav signal, and the non-runaway bubble dynamics model is effectively ruled out by the NANOGrav data at a significance level of $5\sigma$. Secondly, we obtain a remarkably distinct result compared to the previous work, i.e., the bubble dynamics in the phase transition should be characterized by the Jouguet detonation rather than a non-runaway scenario. Thirdly, our results provide some insights into the applicability of holographic descriptions of the quark phase transitions, in which the models of hard and soft walls relate to the introduction of sharp and smooth cutoffs respectively. According to our analysis, the holographic description using the hard wall model for the quark phase transition is more likely to be consistent with PTA data as compared to that using the soft wall model.
In this work, we only examine the first-order confinement-deconfinement phase transitions. However, it is also worth considering other types of the first-order phase transition at the QCD scale, such as the chiral symmetry-breaking Reya (1974). Moreover, relying solely on PTA observations might not be sufficient to distinguish the various details of different first-order phase transitions. Combining additional constraints Bai and Korwar (2022) related to first-order phase transitions will help us gain further insights into the processes of cosmic phase transitions. We leave these aspects for future research.